The Oriole Products Co. currently has debt with a market value of $275 million outstanding. The debt consists of 9 percent coupon bonds (semiannual coupon payments) which have a maturity of 15 years and are currently priced at $1,429.26 per bond. The firm also has an issue of 2 million preferred shares outstanding with a market price of $14 per share. The preferred shares pay an annual dividend of $1.20. Oriole also has 14 million shares of common stock outstanding with a price of $20.00 per share. The firm is expected to pay a $2.20 common dividend one year from today, and that dividend is expected to increase by 4 percent per year forever. If Oriole is subject to a 40 percent marginal tax rate, then what is the firm’s weighted average cost of capital?

Calculate the weights for debt, common equity, and preferred equity. (Round intermediate calculations and final answers to 4 decimal places, e.g. 1.2514.)

Debt
Preferred equity
Common equity

Calculate the cost of debt. (Round intermediate calculations to 4 decimal places, e.g. 1.2514 and final answer to 2 decimal places, e.g. 15.25%.)

Cost of debt

Calculate the cost of preferred equity. (Round intermediate calculations to 4 decimal places, e.g. 1.2514 and final answer to 2 decimal places, e.g. 15.25%.)

Cost of preferred equity

Calculate the cost of common equity. (Round intermediate calculations to 4 decimal places, e.g. 1.2514 and final answer to 0 decimal places, e.g. 15%.)

Cost of common equity

What is the firm’s weighted average cost of capital? (Round intermediate calculations to 4 decimal places, e.g. 1.2514 and final answer to 2 decimal places, e.g. 15.25%.)

WACC
Answer #1

x V 3x =+C165/$C$174 0165 ДА Market After-tax cost % Market value weights 15.00% $ 280,000,000 10.3994 =2.2/20+4% =14000000*2

*Please rate thumbs up